Corporate Finance – Chapter 27- Short Term Financial PlanningQ. 05: The Treadwater Bank wants to raise $1.1 million using three-month commercial paper. The net proceeds to the bank will
Answer #1

Answer 5)

Raised $ OR value to be reapid = 1,100,000

Net proceeds received = 1,084,600

Excess Paid at Maturity = 1,100,000 - 1,084,600 = 15,400

Interest Rate for 3 months = Excess paid / Net Proceeds

= 15,400 / 1,084,600  

= 1.4199%

Effective Rate = (1+ 0.014199 ) ^4 -1

= 1.05801716097 -1

= 5.80%

Answer 6)

Interest per quarter = 488

Principal = 14000

Interest Rate for a quarter = 488 / 14000

= 0.03485714285

Effective Interest Rate = (1+ 0.03485714285) ^4-1

= 1.14699857866 -1

= 14.69%  

Answer 7)

Raised $ OR value to be reapid = 6,300,000

Net proceeds received = 6,113,898

Excess Paid at Maturity = 6,300,000 - 6,113,898 = 186,102

Interest Rate for 3 months = Excess paid / Net Proceeds

= 186,102 / 6,113,898

= 3.043917317%

Effective Rate = (1+ 0.03043917317 ) ^3 -1

= 1.0941253525 -1

= 9.41%